# Some Super-approximation Rates of ReLU Neural Networks for Korobov Functions

### v1

@@ 5 Conclusion and Future Work
Using the bit-extraction technique, we have proved nearly optimal super-approximation $L_{p}$ and $W^{1}_{p}$ error bounds of ReLU DNNs for Korobov functions ($1\leq p<\infty$). Furthermore, we also provide remarks on the approximation performance under certain other network structures and norms. The bit extraction techniques based on ReLU networks, Floor-ReLU networks (Shen et al. (2021a)), or FLES networks (Shen et al. (2021b)) have been well developed. Nevertheless, investigating bit extraction techniques based on other activation functions will lead to a range of network approximation characteristics across various function spaces. This will remain a future research direction.

### v2

@@ 5 Conclusion and Future Work
Using the bit-extraction technique, we have proven nearly optimal super-approximation $L_{p}$ and $W^{1}_{p}$ error bounds of ReLU DNNs for Korobov functions ($1\leq p<\infty$). Furthermore, we also provide remarks on the approximation performance under certain other network structures and norms. The bit extraction techniques based on ReLU networks, Floor-ReLU networks ([32]), or FLES networks ([33]) have been well developed. Nevertheless, investigating bit extraction techniques based on other activation functions will lead to a range of network approximation characteristics across various function spaces. This will remain a future research direction.
Acknowledgments
Y. L. was supported by the National Natural Science Foundation of China under grant 12471346. The authors would like to thank the anonymous referees for helpful remarks that improved the quality and presentation of this paper.